# Dissecting Mahalanobis: How Feature Geometry and Normalization Shape OOD Detection

### v1

@@ Abstract
Out-of-distribution (OOD) detection is critical for the reliable deployment of deep learning models. hile Mahalanobis distance methods are widely used, the impact of representation geometry and normalization on their performance is not fully understood, which may limit their downstream application. To address this gap, we conducted a comprehensive empirical study across diverse image foundation models, datasets, and distance normalization schemes. First, our analysis shows that Mahalanobis-based methods aren’t universally reliable. Second, we define the ideal geometry for data representations and demonstrate that spectral and intrinsic-dimensionality metrics can accurately predict a model’s OOD performance. Finally, we analyze how normalization impacts OOD performance. Building upon these studies, we propose radially scaled $\ell_{2}$ normalization, a method that generalizes the standard $\ell_{2}$ normalization recently applied to Mahalanobis-based OOD detection. Our approach introduces a tunable parameter to directly control the radial geometry of the feature space, systematically contracting or expanding representations to significantly improve OOD detection performance. By bridging the gap between representation geometry, normalization, and OOD performance, our findings offer new insights into the design of more effective and reliable deep learning models.
@@ 1 Introduction
Out-of-distribution (OOD) detection is foundational for building reliable, open-world vision systems, yet consistent evaluation at scale—especially with modern foundation models—remains challenging and essential for practice. Mahalanobis-based detectors (Lee et al., 2018) are surprisingly simple yet powerful baselines that often achieve state-of-the-art performance (Mueller & Hein, 2025; 2024). At its core, this approach models the feature distribution of in-distribution data—typically by fitting class-conditional multivariate Gaussians—and flags an input as OOD if its feature representation is far from all class centroids. While effective, it is not fully understood why this simple metric works so well or how the complex geometry of high-dimensional representations contributes to its success. This paper systematically investigates this question, revealing that representation geometry and feature normalization are the primary drivers of Mahalanobis-based OOD detection performance and providing a practical method to optimize them.
Figure 1: Effect of $\beta$-scaled $\ell_{2}$ normalization effect 2D feature geometry and Mahalanobis decision surfaces. Optimal $\beta$ improve OOD detection performance. Gray arrows indicate the mapping from the original to the transformed space. Increasing $\beta$ contracts norms and tightens clusters, yielding smoother, more localized decision regions; negative $\beta$ spreads points and broadens regions. Choosing an appropriate $\beta$ improves class separation and downstream OOD detection.
We begin our work by benchmarking a diverse set of self-supervised models, revealing significant variance in the inherent OOD detection capabilities of their representations. We then demonstrate that this variance is not random, but correlates strongly with measurable geometric properties of the in-distribution feature space, such as its intrinsic dimensionality and spectral structure. Unlike prior works Ren et al. (2021); Mueller & Hein (2025) focused on refining the distance metric itself, we introduce a direct control knob for the geometry of the representation: a $\beta$-scaled $\ell_{2}$ normalization that radially contracts or expands the feature space. This allows us to reshape the representation to better suit the detector. Finally, we demonstrate that a simple regression model can predict this optimal $\beta$, without access to target OOD data, achieving nearly the same performance as an oracle with access to OOD samples. Our main contributions are:
@@
    A comparative study of several Mahalanobis-based OOD detectors across diverse models, including a per-dimension analysis of its variants.
@@
    An analysis revealing that OOD performance is strongly predicted by the in-distribution geometry of features, such as spectral decay and intrinsic rank.
@@
    The introduction of $\beta$-scaled $\ell_{2}$ normalization and a method to predict the optimal $\beta$ using only in-distribution data.
@@ 2 Related Work
Out-of-distribution (OOD) detection is essential for ensuring the reliability of machine learning systems in real-world deployment (Fort et al., 2021). Its goal is to identify whether inputs stem from the training distribution, thus preventing overconfident predictions on unexpected data (Yang et al., 2024). Post-hoc, training-free methods are particularly effective, as they combine efficiency with robustness without altering the model (Xu et al., 2023). Among OOD detection methods, Mahalanobis distance has become a cornerstone (Lee et al., 2018), with several refinements improving its robustness and performance. The standard Mahalanobis distance (MD) uses class-conditional covariance estimates to measure the distance of a sample from each class mean. In contrast, the Relative Mahalanobis distance (RMD) (Ren et al., 2021) compares each class-specific distance to a single global Gaussian fitted to all in-distribution (ID) data, effectively normalizing class distances against a global reference. Mahalanobis++ (Mueller & Hein, 2025) further improves performance by L2-normalizing features, making them adhere more closely to the Gaussian assumptions underlying the Mahalanobis distance. However, our sudy reveals broader insight on the influence of normalization while computing Mahalanobis distance, particularly in the context of vision models.
@@
Vision OOD detection has shifted toward leveraging large-scale pretraining and contrastive objectives, where vision transformers Dosovitskiy et al. (2021) and CLIP Radford et al. (2021) show strong near-OOD performance and benefit markedly from few-shot outlier exposure and even label-only supervision for outlier classes (Fort et al., 2021). However, full fine-tuning can distort pretrained representations and harm OOD generalization relative to linear probing, with similar cautions for vision–language models; recent work also explores training-time scaling and post-hoc enhancements, and revisits detector design in vision foundation models (Fort et al., 2021; Ming & Li, 2024; Xu et al., 2023; Zhao et al., 2024b). Evaluation rigor has improved through ImageNet-scale suites like NINCO that mitigate in-distribution leakage, while theory and diagnostics connect feature separability to OOD error and delineate when OOD detection is learnable (Bitterwolf et al., 2023; XIE et al., 2023).
@@
Representation geometry and normalization have attracted increased attention for their role in OOD generalization. Analyses of contrastive learning and normalization approaches (Le-Gia & Ahn, 2023; Tan et al., 2025) show that geometric priors, such as hyperspherical projection or $\ell_{2}$ normalization, can yield more robust representation spaces. Studies like (Zhao et al., 2024a) and (XIE et al., 2023) link improved feature separability and lower intrinsic dimensionality to higher OOD detection performance.
@@ 3 Comparative Study of Self-Supervised Models
3.1 Background
Let $z^{\prime}=f(x^{\prime})\in\mathbb{R}^{d}$ be the feature representation of a test image input $x^{\prime}$, and let $\{\mathcal{N}(\mu_{k},\Sigma)\}_{k=1}^{K}$ denote the $K$ class–conditional Gaussian distributions fitted to in–distribution (ID) training data.
Mahalanobis distance (MD) (Lee et al., 2018) measures the squared distance of $z^{\prime}$ from the mean of each class: $\operatorname{MD}_{k}(z^{\prime})\;=\;(z^{\prime}-\mu_{k})^{\top}\Sigma^{-1}(z^{\prime}-\mu_{k}).$ A confidence score is obtained as the negative minimum distance, $\mathcal{C}(x^{\prime})=-\min_{k}\operatorname{MD}_{k}(z^{\prime}).$
Relative Mahalanobis distance (RMD) (Ren et al., 2021) compares the class-specific distance to a single global Gaussian fitted to all ID data, the marginal Mahalanobis distance (MMD): $\operatorname{MD}_{0}(z^{\prime})=(z^{\prime}-\mu_{0})^{\top}\Sigma_{0}^{-1}(z^{\prime}-\mu_{0}),$ and defines the RMD score as: $\operatorname{RMD}_{k}(z^{\prime})=\operatorname{MD}_{k}(z^{\prime})-\operatorname{MD}_{0}(z^{\prime}).$ This effectively normalizes class distances against a global reference.
Eigenvalue Decomposition.
Let $\Sigma=U\Lambda U^{\top}$ be the eigendecomposition of the shared covariance matrix, where $\Lambda=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{d})$ and $U$ is orthonormal. As shown in (Mueller & Hein, 2025), in this basis the regular Mahalanobis distance decomposes as
$\operatorname{MD}_{k}(z^{\prime})=\sum_{i=1}^{d}\frac{\bigl[u_{i}^{\top}(z^{\prime}-\mu_{k})\bigr]^{2}}{\lambda_{i}},$ (1)
revealing the contribution of each principal component. The eigenvalues $\{\lambda_{i}\}$ quantify the spread of in-distribution features along each direction; small $\lambda_{i}$ correspond to directions of small variance and thus high discriminative power for OOD detection.
@@ 3.2 Cross-model OOD Detection Performance
Figure 2: OOD detection performance across model families on the NINCO. RMD consistently outperforms the standard MD, especially for models pretrained but not fine-tuned on ImageNet.
Understanding how different representation learning strategies affect OOD detection is a critical first step in our study. Mahalanobis-based detectors are widely used, but their sensitivity to model architecture, pretraining data, and fine-tuning is not well characterized. We therefore begin with a broad, model-agnostic comparison to answer a simple question: Which modern self-supervised or pretrained vision models produce representations that naturally lend themselves to Mahalanobis-style OOD detection?
To this end, we gathered publicly available checkpoints from timm (Wightman, 2019) and huggingface-transformers (Wolf et al., 2020), spanning a range of architectures, sizes, and pretraining objectives. Following the OpenOOD protocol (Yang et al., 2022), we evaluate on five standard benchmarks—NINCO (Bitterwolf et al., 2023), iNaturalist (Van Horn et al., 2018), SSB-Hard (Bitterwolf et al., 2023), OpenImages-O (Krasin et al., 2017), and Textures (Cimpoi et al., 2014). Performance is reported as the false positive rate at 95% true positive rate (FPR@95 or FPR). Additional implementation details and a full model list appear in the Appendix.
Key Insights (Figure 2)
(1) RMD consistently improves performance over standard Mahalanobis distance, with the largest gains in models pretrained—but not fine-tuned—on ImageNet. For example, RMD markedly boosts the OOD detection of EVA02-In21k and ViT-In21k, matching or surpassing their fine-tuned counterparts. This weakens the usual correlation between in-distribution accuracy and FPR and yields more uniform score distributions. (2) Classification accuracy is not a reliable proxy for OOD performance. Substantial accuracy gaps (often $10\%$) do not necessarily translate into improved detection, though we observe a mild correlation along the fine-tuning sequence In1k $\rightarrow$ In22k-In1k $\rightarrow$ large In22k-In1k models (complete results in Appendix D.1).
@@ 3.3 Mahalanobis Variants and Per-Dimension Analysis
Having established cross-model trends, we next ask: Which aspects of the Mahalanobis representation space actually drive OOD discrimination? Beyond aggregate scores, the structure of individual feature dimensions may reveal why certain models excel while others falter. We therefore conduct a detailed per-dimension investigation using three Mahalanobis variants—regular, marginal, and relative. Using the decomposition in Eq. equation 1, we define the OOD separation of the $i$-th eigenvector direction as the difference between its mean contribution for out-of-distribution samples and for in-distribution samples: $S_{i}=\mathbb{E}_{x^{\prime}\sim\mathcal{D}_{\text{OOD}}}\!\left[\frac{\bigl[u_{i}^{\top}(z^{\prime}-\mu_{k})\bigr]^{2}}{\lambda_{i}}\right]-\mathbb{E}_{x\sim\mathcal{D}_{\text{ID}}}\!\left[\frac{\bigl[u_{i}^{\top}(z-\mu_{k})\bigr]^{2}}{\lambda_{i}}\right].$ Here $S_{i}$ quantifies how strongly dimension $i$ separates OOD from in-distribution data; positive values indicate greater OOD spread along that eigen-direction. The eigenvalues $\{\lambda_{i}\}$ are sorted in descending order ($\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{d}$), so dimension $i$ on subsequent plots corresponds to the $i$-th largest eigenvalue.
Figure 3 presents two complementary analyses. The top row reports OOD separation, exposing which latent directions contribute most to detection. The bottom row shows a dimension-ablation study: we incrementally compute FPR using the first $K$ principal components (forward) or start from the least-variant dimensions (backward).
Figure 3: Dimension-wise analysis of OOD separation (top) and FPR under progressive dimension ablation (bottom). Large embedding-space separation does not necessarily guarantee superior detection.
Key Insights
(1) Large OOD separation $S_{i}$ do not always yield lower FPR: for example, BEiTV2 FT In1k exhibits stronger separation across all three distance metrics yet performs on par with—or worse than—BEiTV2 FT In21k. (2) The number of dimensions required for optimal detection varies widely: some models saturate quickly, whereas others need nearly the full spectrum. Backward ablation reveals that ViT In21k achieves its best FPR using only the second half of the spectrum, indicating that directions of smaller explained variance can be more discriminative for OOD-ness. (3) MMD results show that certain models rely heavily on class-discriminative features, while others—such as CLIP—spread OOD samples far from ID data regardless of their classes, consistently achieving low MMD scores.

### v3

# A Geometry-Based View of Mahalanobis OOD Detection

@@ Abstract
Out-of-distribution (OOD) detection is critical for reliable deployment of vision models. Mahalanobis-based detectors remain strong baselines, yet their performance varies widely across modern pretrained representations, and it is unclear which properties of a feature space cause these methods to succeed or fail. We conduct a large-scale study across diverse foundation-model backbones and Mahalanobis variants. First, we show that Mahalanobis-style OOD detection is not universally reliable: performance is highly representation-dependent and can shift substantially with pretraining data and fine-tuning regimes. Second, we link this variability to in-distribution geometry and identify a two-term ID summary that consistently tracks Mahalanobis OOD behavior across detectors: within-class spectral structure and local intrinsic dimensionality. Finally, we treat normalization as a geometric control mechanism and introduce radially scaled $\ell_{2}$ normalization, $\phi_{\beta}(z)=z/\|z\|^{\beta}$, which preserves directions while contracting or expanding feature radii. Varying $\beta$ changes the radii while preserving directions, so the same quadratic detector sees a different ID geometry. We choose $\beta$ from ID-only geometry signals and typically outperform fixed normalization baselines.
Machine Learning, Out-of-Distribution Detection, Mahalanobis Distance
@@ 1 Introduction
Figure 1: Effect of radially scaled $\ell_{2}$ normalization on feature geometry and Mahalanobis boundaries. Visualization in 2D of the induced geometry and the resulting Mahalanobis decision regions. Gray arrows indicate the radial mapping from the original to the transformed space. Larger $\beta$ contracts feature norms and tightens clusters, producing more localized decision regions; smaller $\beta$ expands norms and broadens them. Choosing $\beta$ appropriately can reduce ID–OOD overlap and improve OOD detection.
Mahalanobis scores are among the simplest post-hoc detectors for out-of-distribution (OOD) detection, yet they remain surprisingly competitive on modern vision backbones (Lee et al., 2018; Mueller & Hein, 2025, 2024). At the same time, their behavior is highly representation-dependent: the same quadratic detector can perform well on one pretrained model and fail on another, and performance can shift sharply with changes in pretraining data or fine-tuning regime. This sensitivity makes Mahalanobis-based OOD detection difficult to deploy reliably and raises a basic question: which properties of an in-distribution feature space determine when a Mahalanobis detector succeeds or fails?
We study Mahalanobis OOD detection through the lens of representation geometry. Across self supervised and foundation model representations, we show that geometric structure accounts for much of the observed cross model variation. In particular, a compact ID-only summary combining local intrinsic dimensionality and within-class spectral decay strongly predicts Mahalanobis OOD performance across variants. This connects detector reliability to measurable properties of the in-distribution feature space.
Motivated by this geometric view, we introduce a simple post-hoc control mechanism that changes the geometry presented to the same quadratic detector. We use radially scaled $\ell_{2}$ normalization, $\phi_{\beta}(z)=z/\|z\|^{\beta}$, which preserves feature directions while contracting or expanding radii. Unlike prior work that modifies the scoring rule (Ren et al., 2021) or fixes normalization to the unit sphere (Mueller & Hein, 2025), varying $\beta$ provides a continuous way to reshape radial geometry without altering the detector form. Figure 1 illustrates how $\beta$ tightens or spreads decision regions by changing feature radii. Empirically, adjusting $\beta$ induces structured, model-specific changes in both geometry and OOD performance. Leveraging the same ID-only geometry signals, we propose a practical procedure to select $\beta$ without access to OOD samples, often improving over fixed baselines such as $\beta=0$ (standard features) and $\beta=1$ (unit-sphere normalization).
Our main contributions are:
@@
    A broad benchmark of Mahalanobis-style OOD detectors across diverse SSL/foundation models, including a per-dimension analysis of detector behavior.
@@
    An empirical link between Mahalanobis OOD performance and ID geometry, including an ID-only summary that consistently predicts performance across detector variants.
@@
    A geometric control mechanism via $\beta$-scaled radial normalization, together with an ID-only $\beta$ selection rule that approaches oracle tuning without requiring OOD access.
@@ 2 Related Work
OOD detection is essential for ensuring the reliability of machine learning systems in real-world deployment (Fort et al., 2021). Its goal is to identify whether inputs stem from the training distribution, thus preventing overconfident predictions on unexpected data (Yang et al., 2024). Post-hoc, training-free methods are particularly effective, as they combine efficiency with robustness without altering the model (Xu et al., 2023). Among OOD detection methods, Mahalanobis distance has become a cornerstone (Lee et al., 2018), with several refinements improving its robustness and performance. The standard Mahalanobis distance (MD) uses class-conditional covariance estimates to measure the distance of a sample from each class mean. In contrast, the Relative Mahalanobis distance (RMD) (Ren et al., 2021) compares each class-specific distance to a single global Gaussian fitted to all in-distribution (ID) data, effectively normalizing class distances against a global reference. Mahalanobis++ (Mueller & Hein, 2025) further improves performance by L2-normalizing features, making them adhere more closely to the Gaussian assumptions underlying the Mahalanobis distance. However, our study reveals broader insight into the influence of normalization when computing Mahalanobis distance, particularly in the context of vision models.
@@
Vision OOD detection has shifted toward leveraging large-scale pretraining and contrastive objectives, where vision transformers (Dosovitskiy et al., 2021) and CLIP (Radford et al., 2021) show strong near-OOD performance and benefit markedly from few-shot outlier exposure and even label-only supervision for outlier classes (Fort et al., 2021). However, full fine-tuning can distort pretrained representations and harm OOD generalization relative to linear probing, with similar cautions for vision–language models; recent work also explores training-time scaling and post-hoc enhancements, and revisits detector design in vision foundation models (Fort et al., 2021; Ming & Li, 2024; Xu et al., 2023; Zhao et al., 2024b). Evaluation rigor has improved through ImageNet-scale suites like NINCO that mitigate in-distribution leakage. Meanwhile, theory and diagnostics connect feature separability to OOD error and delineate when OOD detection is learnable (Bitterwolf et al., 2023; Xie et al., 2023).
@@
Representation geometry and normalization have attracted increased attention for their role in OOD generalization. An analysis of contrastive learning and normalization approaches (Le-Gia & Ahn, 2023; Tan et al., 2025) shows that geometric priors, such as hyperspherical projection or $\ell_{2}$ normalization, can yield more robust representation spaces. Studies like (Zhao et al., 2024a) and (Xie et al., 2023) link improved feature separability and lower intrinsic dimensionality to higher OOD detection performance.
Figure 2: Dimension-wise OOD behavior for Mahalanobis variants. The top row shows dimension-wise OOD separation $S_{i}$, and the bottom row reports FPR under progressive dimension ablation. Strong embedding-space separation does not necessarily translate to better OOD detection.
3 Background: Mahalanobis OOD Detection
Let $z=f(x)\in\mathbb{R}^{d}$ denote the feature representation of an input $x$. Given ID training features from $K$ classes, Mahalanobis-based detectors model each class by a Gaussian $\mathcal{N}(\mu_{k},\Sigma)$ with a tied covariance $\Sigma$ (the LDA assumption), and optionally a marginal Gaussian $\mathcal{N}(\mu_{0},\Sigma_{0})$ fitted to all ID features.
Mahalanobis variants.
The class-conditional Mahalanobis distance (MD) and its confidence score are
$\operatorname{MD}_{k}(z)$ $=(z-\mu_{k})^{\top}\Sigma^{-1}(z-\mu_{k}),$ (1)
$\mathcal{C}_{\mathrm{MD}}(x)$ $=-\min_{k}\operatorname{MD}_{k}\!\big(f(x)\big).$
Marginal Mahalanobis (MMD) uses the class-agnostic quadratic form $\operatorname{MD}_{0}(z)=(z-\mu_{0})^{\top}\Sigma_{0}^{-1}(z-\mu_{0})$ as its score, and Relative Mahalanobis (RMD) (Ren et al., 2021) subtracts this marginal reference:
$\operatorname{RMD}_{k}(z)$ $=\operatorname{MD}_{k}(z)-\operatorname{MD}_{0}(z),$ (2)
$\mathcal{C}_{\mathrm{RMD}}(x)$ $=-\min_{k}\operatorname{RMD}_{k}(z).$
Eigenbasis view.
Let $\Sigma=U\Lambda U^{\top}$ have eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{d}>0$ and eigenvectors $\{u_{i}\}$. Define the per-direction energy $\tilde{a}_{i}(z)\triangleq(u_{i}^{\top}(z-\mu_{k}))^{2}$. Then
$\operatorname{MD}_{k}(z)=\sum_{i=1}^{d}\lambda_{i}^{-1}\,\tilde{a}_{i}(z).$ (3)
This form highlights how directions with smaller variance (smaller $\lambda_{i}$) receive larger inverse weighting and motivates our later analysis of how representation geometry influences Mahalanobis-style OOD behavior.
@@ 4.1 Cross-model OOD Detection Performance
We begin by characterizing how representation learning choices shape Mahalanobis-style OOD detection across modern vision backbones. In particular, we ask how performance depends on architecture, pretraining data, and fine-tuning regime, i.e., factors whose effects are not systematically documented. This motivates a broad, model-agnostic comparison: Which modern self-supervised or pretrained vision models produce representations that naturally lend themselves to Mahalanobis-style OOD detection?
Models.
We evaluate publicly available checkpoints from timm (Wightman, 2019) and huggingface-transformers (Wolf et al., 2020), covering multiple transformer families, model scales, and training objectives. The full model list is provided in Appendix I.
Evaluation protocol.
Following OpenOOD (Yang et al., 2022), ImageNet-1K serves as the in-distribution (ID) dataset (train features for fitting; validation for ID testing). We report $\mathrm{FPR}@95$ for distinguishing ImageNet validation from each of the five OOD benchmarks: NINCO (Bitterwolf et al., 2023), iNaturalist (Van Horn et al., 2018), SSB-Hard (Bitterwolf et al., 2023), OpenImages-O (Krasin et al., 2017), and Textures (Cimpoi et al., 2014). Unless stated otherwise, we fit class means $\{\mu_{k}\}$ and a tied covariance $\Sigma$ on ImageNet-1K training features and evaluate OOD scores on the ImageNet validation set versus each OOD dataset.
Results (Figure 3).
RMD improves over the standard Mahalanobis distance in most settings, with the largest gains for models that are pretrained but not fine-tuned on ImageNet. Notably, RMD substantially improves OOD detection for EVA02-In21k and ViT-In21k, in some cases matching or exceeding the performance of their ImageNet-fine-tuned counterparts. This weakens the typical association between in-distribution accuracy and FPR, and produces more consistent score distributions across models. At the same time, classification accuracy is not a reliable proxy for OOD performance: large accuracy gaps (often $10\%$) do not necessarily yield better detection. We observe only a mild trend along the fine-tuning sequence In1k $\rightarrow$ In22k-In1k $\rightarrow$ larger In22k-In1k models; full results are reported in Appendix F.1.
Figure 3: OOD detection performance on NINCO across model families. RMDS consistently improves over standard MD, with the largest gains for models that are pretrained but not fine-tuned on ImageNet.
Figure 4: Correlations between representation metrics and OOD performance. Spearman correlations between representation metrics and OOD performance across Mahalanobis variants. Here, F denotes metrics computed directly from feature representations, whereas C, $\mathbf{S_{b}}$, and $\mathbf{S_{w}}$ are derived from the eigenvalue spectra of the covariance/scatter matrices (Appendix A). The three Mahalanobis-based detectors leverage distinct geometric cues, yielding different correlation patterns. Similar trends hold for Pearson correlations (Appendix H).
@@ 4.2 Mahalanobis Variants and Per-Dimension Analysis
Having established cross-model trends, we next ask: which parts of the representation space drive OOD discrimination? Beyond aggregate scores, inspecting individual eigen-directions helps explain why some models detect OOD reliably while others do not. We therefore analyze MD, MMD, and RMD at the level of their per-direction contributions.
Per-direction separation.
Using the decomposition in Eq. 3, we define the OOD separation in the eigen-direction $i$ as the difference between its mean contribution to OOD and ID:
$S_{i}\;\triangleq\;\lambda_{i}^{-1}\!\left(\mathbb{E}_{x\sim\mathcal{D}_{\mathrm{OOD}}}[\tilde{a}_{i}(x)]-\mathbb{E}_{x\sim\mathcal{D}_{\mathrm{ID}}}[\tilde{a}_{i}(x)]\right),$ (4)
where the projection term $\tilde{a}_{i}(\cdot)$ utilizes the selected class $k\in\arg\min_{c}\operatorname{MD}_{c}$ (or $\arg\min_{c}\operatorname{RMD}_{c}$ for RMD; see Eq. 1 and 2). $S_{i}$ measures the difference between average ID and OOD Mahalanobis energy along eigen-direction $i$. Positive $S_{i}$ indicates that OOD samples contribute more inverse-variance weighted energy than ID samples along $u_{i}$. We order eigenvalues as $\lambda_{1}\geq\cdots\geq\lambda_{d}$, so larger $i$ corresponds to smaller-variance directions.
Ablation protocol.
Figure 2 (bottom) reports an ablation study where we recompute the FPR using only the first $K$ eigen-directions (forward ablation) or only the last $K$ directions (backward ablation). This isolates whether discrimination arises from high-variance structure (small $i$) or from low-variance components that receive strong inverse-variance weighting (large $i$).
Results.
Figure 2 reveals three recurring behaviors across models. (i) Large per-direction separation does not necessarily yield low FPR: for example, BEiTV2 FT In1k exhibits stronger separation across many directions yet performs on par with (or worse than) BEiTV2 FT In21k. (ii) The effective number of eigen-directions needed for strong detection varies substantially: some models saturate quickly with small $K$, while others require most of the spectrum to approach their best FPR. (iii) Backward ablation shows that low-variance directions can dominate discrimination in some settings; e.g., ViT In21k attains its best FPR primarily from the latter part of the spectrum, suggesting that small-variance components can carry disproportionate OOD signal after inverse-variance weighting. These observations motivate our later stability analysis: performance is governed not only by mean ID–OOD separation but also by how quadratic weighting interacts with the representation spectrum and the allocation of sample energy across eigen-directions.